OpenStudy (anonymous):

Which shows this equation in standard form? 7 – 3(x – y) = 5x + 2 A.3x – 8y = 5 B.8x + 3y = 5 C.8x – 3y = 5 D.8x – 3y = –5 @iGreen

OpenStudy (compassionate):

The Standard Form equation of a line has the following formula: Ax + By = C where A ≠0 and B ≠0

OpenStudy (igreen):

Just simplify. Distribute -3 into the parenthesis: 7 - 3(x - y) = 5x + 2 7 - 3x + 3y = 5x + 2 Subtract 7 to both sides: -3x + 3y = 5x - 5 Add 5x to both sides. -3x + 5x + 3y = -5 Can you add -3x + 5x? (It's just like adding -3 + 5, but you put an 'x' after it.)

OpenStudy (igreen):

@NunChuckBoy3101

OpenStudy (igreen):

No, -3 + 5 does not equal 8..

OpenStudy (igreen):

No, -3 + 5 is the same thing as 5 - 3, what's 5 - 3?

OpenStudy (anonymous):

2x @iGreen

OpenStudy (igreen):

Yep, you got it.

OpenStudy (igreen):

Argh, I made a mistake..you have to subtract 5x to both sides. So what's -3 - 5?

OpenStudy (anonymous):

@iGreen 2x